OpenStudy (anonymous):

can you help me solve this question (20x-10)+(10x+12)=180

OpenStudy (anonymous):

This can be rearranged to 20x + 10x - 10 + 12 = 180. 20x+10x = 30x, -10 +12 = 2. Therefore 30x + 2 = 180. 30x = 178. Which comes to be about x = 89/15 or = 5.93333

OpenStudy (anonymous):

thank you for the help

OpenStudy (anonymous):